### v2

# Confidence is detection-like in high-dimensional spaces

Abstract
Confidence estimates are often “detection-like”—driven by positive evidence in favour of a decision. This empirical observation has been interpreted as showing human metacognition is limited by biases or heuristics. Here we show that Bayesian confidence estimates also exhibit heightened sensitivity to decision-congruent evidence in higher-dimensional signal detection theoretic spaces, leading to detection-like confidence criteria. This effect is due to a nonlinearity induced by normalisation of confidence by a large number of unchosen alternatives. Our analysis suggests that detection-like confidence is rational when participants consider a greater number of hypotheses than assumed by the experimenter.
Keywords metacognition; confidence; inference; positive evidence bias; signal detection theory
1 Introduction
Understanding how confidence is formed in human decision-making is important for identifying sources of metacognitive failure. One prominent metacognitive bias is termed the “positive evidence bias” (PEB). A PEB is the empirical observation that confidence ratings (but not first-order decisions) are often relatively insensitive to evidence against one’s choice [1, 2, 3, 4, 5]. A consequence of a PEB is that observers may become inappropriately confident when the total evidence supporting any decision option goes up, even if their choice accuracy remains unaltered [6].
A PEB can be accounted for within signal detection theory (SDT) as confidence criteria becoming “detection-like” [7, 8, 9]. For a discrimination decision between two symmetrical hypotheses, the total evidence should be irrelevant to one’s confidence in the decision – instead, what matters is the balance of evidence supporting one or other decision option (Figure 1.). For a two-dimensional SDT problem, this yields optimal decision and confidence criteria that track relative evidence (the distance to the main diagonal), irrespective of the total evidence (the distance along the main diagonal). In a detection task, in contrast, the optimal decision and confidence criteria become perpendicular to the evidence axes – because what matters for detecting (rather than discriminating) a stimulus is the extent to which it deviates from the distribution of percepts expected under the absence of a stimulus (the lower-left corner of the space).
Figure 1: Discrimination and detection criteria in a 2-dimensional signal detection theory (SDT) model. Circles indicate bivariate Gaussian stimulus distributions. Reproduced from Mazor et al. [8].
Previous efforts to explain a PEB have focused on explaining why confidence criteria become detection-like. These explanations include proposals that confidence estimates are based on response-congruent evidence [10, 11]; that neural computations of confidence are biased towards accessing a subset of less normalised, “detection-like” neurons [12]; or that there is an additional contribution of stimulus visibility to confidence estimates [7]. Other studies have suggested that a PEB may arise due to a violation in the SDT assumption of equal-variance bivariate Gaussian distributions with naturalistic stimuli, which tend to show more variance along target vs. non-target dimensions [11]. More broadly, the PEB has typically been considered a bias or suboptimality that plagues human metacognition [13, 10, 11].
Against this backdrop, Webb et al. [14] showed that a PEB can emerge as a natural consequence of convolutional neural networks trained to optimise the accuracy of both choices and confidence estimates when discriminating between handwritten digits. These intriguing results indicate a PEB may emerge as a result of the adaptive consequences of learning the statistics of natural stimuli, rather than as a suboptimality peculiar to human metacognition.
Webb et al.’s findings are important in exposing the possibility that apparent suboptimalities in human confidence may be explained by optimal computations applied to more naturalistic evidence spaces. Their results were obtained as a by-product of training a neural network on higher-dimensional stimulus sets. However, if this hypothesis is correct, it should be possible to expose such decision rules directly within the simpler framework of signal detection theory (SDT), rather than relying on a neural network to discover their emergence. In this theoretical note, we document how confidence behaves when extending SDT into higher dimensions.
Dimensionality here refers to the number of possible features tracked by the observer. Classical SDT has one feature dimension encoding properties such as brightness or familiarity. 2D SDT allows two feature dimensions which are typically treated as encoding “evidence” for each of two distinct stimuli – such as a right-tilted or left-tilted grating [15, 9]. But even when tasks force participants to classify stimuli into one of two categories defined by the experimenter, it is often natural for participants to tentatively consider a wide range of hypotheses about the stimulus before making a response. For instance, in an orientation discrimination task, observers are unlikely to restrict their perceptual inference to whether the orientation was left or right tilted, but may also by implicitly sensitive to other incidental properties of the stimulus such as phase, spatial frequency, contrast, size, duration, and so on. To account for this possibility, here we allow for greater numbers of feature dimensions such that the evidence space becomes $k$-dimensional. We simulate a typical PEB experiment where subjects are asked to make a 2-choice discrimination between two stimulus categories, $s_{1}$ and $s_{2}$. We then compute the impact of additional “distractor” stimulus dimensions representing evidence for alternative stimulus categories $s_{3:k}$ on the optimal confidence criteria for two-choice discrimination. To preempt our results, we show that confidence criteria are discrimination-like when the hypothesis space is low-dimensional, but become detection-like when the space is higher-dimensional. The appearance of detection-like confidence criteria in higher-dimensional SDT spaces indicates that a PEB can be rational if subjects are computing confidence in a higher-dimensional evidence space that that assumed by the experimenter.
2 Methods
2.1 Signal detection theory in higher-dimensional spaces
We start with a standard 2-dimensional bivariate Gaussian SDT model (Figure 1), and extend this into higher dimensions, with dimensionality indicated by $k$. Each time we add a feature dimension, we add a new orthogonal stimulus distribution with mean $\mu$. For a given dimensionality $k$, sensory evidence $X$ is drawn from the following distribution:
$X\sim N(\mu_{s},\Sigma)$
$\mu_{s}\in\{0,\mu_{target}\}^{k}$
$\Sigma=I_{k}$
where $\mu_{s}$ is a $k$-dimensional vector with $\mu_{target}$ on the $s$th coordinate and 0s elsewhere, $\mu_{target}$ is the observers’ sensitivity to the target dimension, and $I$ indicates the identity matrix. Here we will focus on the case where $\mu_{target}=1$, but similar results obtain for other values of $\mu_{target}$.
Figure 2: A-D) Impact of computing confidence in higher-dimensional SDT spaces for a 2D decision between $s_{1}$ and $s_{2}$ on A) confidence, B) confidence’s sensitivity to evidence for the chosen alternative, C) confidence’s sensitivity to evidence for the unchosen alternative, and D-E) the positive evidence bias (PEB), here computed as the log ratio of the two previous quantities. Confidence is increasingly detection-like (cf. Figure 1) as dimensionality increases, exhibiting increased sensitivity to evidence for the chosen alternative relative to the unchosen alternative. F) Regression coefficients relating chosen and unchosen evidence samples to accuracy (left panel) and confidence (right panel). G) Posterior probabilities of each stimulus class (colored lines) and confidence (black line) along the slice indicated by the dashed white line in panel (A) for $k=10$.
We assume the observer infers the stimulus category by computing the posterior over all possible categories, $P(S|X)$. Then, the observer chooses the available action which maximizes this probability ($a=\operatorname*{argmax}_{i\in\{1,2\}}P(s_{i}|X)$) and rates confidence as the posterior probability of the chosen category ($confidence=P(s_{a}|X)$). The posterior probabilities can be computed using Bayes’ rule:
$confidence=P(s_{a}|X)=\frac{P(s_{a})p(X|s_{a})}{P(s_{1})p(X|s_{1})+P(s_{2})p(X|s_{2})+\ldots+P(s_{k})p(X|s_{k})}$ (1)
Assuming a flat prior across dimensions we obtain:
$confidence=P(s_{a}|X)=\frac{p(X|s_{a})}{p(X|s_{1})+p(X|s_{2})+\ldots+p(X|s_{k})}$ (2)
As we show in the Appendix A, this formula can be solved to show that confidence can be efficiently computed as the softmax of the evidence with inverse temperature equal to the observer’s sensitivity ($\beta=\mu_{target}$):
$confidence=P(s_{a}|X)=\sigma_{\mu_{target}}(X)_{a}=\frac{1}{e^{\mu_{target}(X_{1}-X_{a})}+e^{\mu_{target}(X_{2}-X_{a})}+\ldots+e^{\mu_{target}(X_{k}-X_{a})}}$ (3)
This equation emphasizes that, as $k$ increases, confidence is subject to increasing normalisation. To plot out example surfaces in Figure 2, we set the unchosen evidence values to 0, but similar results are obtained when this is allowed to vary. All simulation code is available via GitHub at https://github.com/smfleming/highDimSDT.
2.2 Neural network models
To ask whether similar effects of dimensionality on confidence formation would obtain in more realistic stimulus sets, we retrained Webb et al.’s convolutional neural networks optimised to predict choice and confidence on subsets of digits drawn from the MNIST dataset (https://github.com/taylorwwebb/performance_optimized_NN_confidence). The classification layer was trained to label the class of the image, and the confidence layer was trained to output a target of 1 if the classification response was correct, and 0 if the classification response was incorrect.
The model architecture was the same as in [14], and involved three major components. The first is a deep neural network encoder $f$ that takes an image $x$, of class $y$. The output of this encoder is then passed to two output layers $g_{class}$ and $g_{conf}$. The layer $g_{class}$ outputs a predicted class for the image and the layer $g_{conf}$ predicts $p(\hat{y}=y)$, the probability that the classification response is correct. $f$ consisted of 3 convolutional layers followed by 3 fully-connected (FC) layers. The convolutional layers had 32 channels each, a kernel size of 3, a stride of 2, batch normalisation, and leaky ReLU nonlinearities with a negative slope of 0.01. The first 2 FC layers had 256 and 128 units, batch normalisation, and leaky ReLU nonlinearities with a negative slope of 0.01. The output of the encoder was generated by the final FC layer, which had 100 units and no nonlinearity or normalisation.
[…]
3 Results
3.1 Discrimination confidence criteria in higher dimensional spaces
As shown in Figure 2A, as the dimensionality of the evidence space increases, confidence surfaces for a two-choice decision becomes more detection-like, and more consistent with a PEB. This effect is due to the presence of additional “distractor” dimensions influencing confidence formation – if these dimensions could be ignored (more formally, if normalisation of confidence could be restricted to the 2D space), then we would continue to observe discrimination-like criteria even at higher dimensionalities.
We can quantify this by computing the absolute change in confidence with respect to the evidence for the chosen stimulus category $a$ and an unchosen stimulus category $i$, and taking the log ratio of the two partial derivatives. When this value is equal to 0, confidence is equally sensitive to changes in the chosen evidence $X_{a}$ and the unchosen evidence $X_{i}$. When this value is greater than 0, confidence is more sensitive to the chosen evidence $X_{a}$ relative to the unchosen evidence $X_{i}$, which is the hallmark of the positive evidence bias. In Appendix B, we show that the positive evidence bias takes the following form:
$\textrm{PEB}_{i}=\textrm{log}\!\left\lvert\frac{\frac{\partial}{\partial X_{a}}P(s_{a}|X)}{\frac{\partial}{\partial X_{i}}P(s_{a}|X)}\right\rvert=\textrm{log}\!\left(1+\sum_{j\notin\{a,i\}}e^{\mu_{target}(X_{j}-X_{i})}\right)$ (4)
The strength of the PEB determined by Equation 4 is plotted in Figure 2D. Since there is only one unchosen alternative for $k=2$, there are no terms in the sum, and so evidence for chosen and unchosen alternatives exhibit equal influences on confidence in this case ($\textrm{PEB}_{i}=0$). The sum gains an additional positive term for every increase in $k$, meaning that the PEB increases without bound as the number of distractor dimensions increases (Figure 2E). Importantly, this holds regardless of which choice the agent makes, regardless of which unchosen dimension we choose to investigate, and regardless of the actual evidence $X$.
We next show that this effect of increasing dimensionality is sufficient to reproduce a classical PEB effect in simulation. Under a range of dimensionalities, we simulate evidence being drawn from $s_{1}$ or $s_{2}$, and record the model’s choice and confidence. We then regressed the chosen and unchosen evidence against both choice and confidence (using a logistic regression for choice, and a linear regression for confidence). As expected, choice was equally and oppositely affected by evidence for and against the chosen option (Figure 2F). In contrast, confidence is driven to a greater extent by the chosen evidence, and this asymmetry increases as dimensionality increases.
3.2 Explaining the influence of increasing dimensionality
Why does this effect obtain? By taking a slice through the space at a constant, fixed value of $X_{2}$, and plotting the posterior for each dimension as a function of $X_{1}$, we can develop an intuition for this phenomenon (Figure 2G). When $X_{1}$ is low (when it is likely to be unchosen), the other distractor dimensions strongly suppress its influence on confidence, leading to a weak influence of unchosen evidence. But when $X_{1}$ is high (when it is likely to be chosen), it outcompetes the distractor dimensions, and confidence steeply rises with chosen evidence. This is due to the effect of an increasing number of distractor dimensions in the feature space entering into the normalisation term in equation 3 as dimensionality ($k$) increases. In other words, a PEB naturally arises from a non-linearity in the function relating confidence to chosen evidence. This is solely due to the normalisation induced by the other feature dimensions in equation 3, rather than any bias or suboptimality in confidence formation.
3.3 Revisiting variance asymmetries
Miyoshi & Lau [11] proposed that a PEB may result from the evidence distributions in SDT having unequal variance along the target and non-target dimensions, leading to the adoption of curved, detection-like confidence criteria. However, their proposal does not explain why an inequality in variance may ensue in the first place. The current simulations in higher-dimensional spaces allow us to revisit this issue and obtain insight into one potential source of the variance inequality.
In Figure 3A, we simulate samples from two equal variance stimulus distributions ($s_{1}$ and $s_{2}$), colour-coded by the decision and level of confidence arising from each evidence sample. In standard 2D SDT, all of the variance in confidence is on the minor diagonal, as expected, and the variances along the target and non-target dimensions are similar. But as dimensionality increases, the variance in confidence becomes greater in the target dimension, and weaker in the non-target dimension (this is a direct consequence of the confidence criteria becoming more detection-like, and parallel to the evidence axes). Importantly, however, this increasing variance in target (vs. non-target) confidence is obtained despite the evidence distributions remaining equal in variance (the two clouds of points remain similar in shape and size across dimensionalities). This can be appreciated by computing the ratio of the variances along the major and minor axes as a function of dimensionality, computed either from the evidence samples (Figure 3B) or the confidence samples (Figure 3C). Only for the latter do we see the variance inequality increasing with dimensionality. In other words, the asymmetry in variance here is in the space of confidence, not in the space of evidence. We will revisit this observation in the Discussion, as this observation leads to an intriguing empirical prediction that it should be possible to “flip” a PEB (in confidence space) simply by changing the target dimension.
Figure 3: A) Samples from two equal variance stimulus distributions ($s_{1}$ and $s_{2}$), colour-coded by the level of confidence attached to a decision arising from each evidence sample, for increasing dimensionality. For a decision of $s_{1}$, confidence increases from blue to green. For a decision of $s_{2}$, confidence increases from red to yellow. B, C) Target:non-target variance ratio computed in (B) evidence space and (C) confidence space.
3.4 Beyond the two-choice case
So far we have considered a two-choice “task” faced by an observer using a higher-dimensional stimulus space in which to compute their (discrimination) confidence. But there is nothing special about the two-choice case. While it is harder to visualise, we can also demonstrate a similar effect in a 3-choice (or N-choice) task [16], where dimensionality $k\geq N$. Figure 4 shows the confidence surfaces for a 3-choice discrimination decision with increasing dimensionality of the evidence space. As for the 2-choice case, when $k=N$, the confidence criteria are aligned along the diagonal planes that separate one stimulus distribution from another. But as $k>N$, the confidence criteria become increasingly parallel to the sides of the cube – becoming increasingly detection-like.
Figure 4: Impact of computing confidence in higher-dimensional SDT spaces on confidence surfaces for a 3-way decision between $s_{1}$, $s_{2}$ and $s_{3}$. Conventions as in Figure 2.
3.5 Effect of dimensionality on confidence formation in neural networks
To ask whether similar effects of dimensionality on confidence formation would obtain in more realistic stimulus sets, we trained convolutional neural networks to predict the confidence and accuracy of MNIST digit classifications [14] (Figure 5). Following Webb et al., we created sets of HPE and LPE stimuli by simultaneously varying the signal and noise within each digit image (Figure 5). We varied the dimensionality of the stimulus space by varying how many (randomly subsampled) digits from 1-9 would be included in the network’s training set (such that dimensionality $k$ ranged from 2 to 10).
[…]
We also considered that the influence of dimensionality on the PEB may be better accommodated by a logarithmic, rather than linear relationship, due to the logarithmic form of Equation 4 depicted in Figure 2E. This was indeed the case, with a logarithmic model providing a better fit as quantified with BIC scores (logarithmic, BIC = -6106.5; linear, BIC = -6092.5). The fits of the logarithmic model are included in Figure 5C.
Figure 5: A) Model architecture (modified from [14]). An image $x$, belonging to class $y$, was passed through a deep neural network (DNN) encoder $f$, followed by two output layers: $g_{class}$ generated a decision classifying the image, and $g_{conf}$ generated a confidence score by predicting $p(\hat{y}=y)$, the probability that the decision was correct. B, C) Impact of varying the dimensionality of the training set on a positive evidence bias in confidence obtained for MNIST digit classification. The fitted lines are derived from a logarithmic model in which the impact of dimensionality on PEB saturates at higher dimensionalities.
3.6 Dimensionality and metacognitive sensitivity
We have shown that if participants consider a richer hypothesis space than assumed by the experimenter, the functional form of the optimal confidence criteria shifts, resulting in a positive evidence bias. We next asked whether dimensionality also affected metacognitive sensitivity—the extent to which confidence discriminates between correct and incorrect responses [17]. In Figure 6A-B, we plot distributions of posterior probabilities and confidence from simulated signal detection agents varying in their number of dimensions. As can be seen in Figure 6A, although posterior probabilities always tend to be larger for target relative to non-target stimuli, these differences become smaller as $k$ increases. Likewise, in Figure 6B, distributions of confidence for correct and incorrect decisions become more and more similar with increasing $k$. Figure 6C plots the corresponding Type 2 ROCs for each value of $k$, which are defined by the Type 2 hit rate (i.e., the proportion of confidence ratings following a correct decision greater than $c$) and the Type 2 false alarm rate (i.e., the proportion of confidence ratings following an incorrect decision greater than $c$) across a range of confidence criteria $c$. Figure 6C shows that the above pattern results in reduced metacognitive sensitivity, meaning that it is increasingly difficult to discriminate between correct and incorrect decisions in high-dimensional hypothesis spaces. Since dimensionality reduces metacognitive sensitivity but has no effect on Type 1 decision accuracy, it follows that dimensionality also reduces metacognitive efficiency (the ratio of metacognitive sensitivity to Type 1 sensitivity) in a similar matter.
Figure 6: Influence of distractor dimensions on (A) the posterior probability of target and non-target stimuli, (B) confidence for correct and incorrect decisions, and (C) the type 2 ROC. For all values of $k$, posterior probabilities tend to be higher for target relative to non-target stimuli, confidence tends to be higher after a correct decision, and meta-d’ is above 0. However, as $k$ increases, differences between target and non-target stimuli become less pronounced, resulting in more similar distributions of confidence for correct and incorrect decisions, and reduced metacognitive sensitivity.
4 Discussion
Here we show that extending SDT into higher dimensional evidence spaces naturally leads to detection-like confidence surfaces when computing confidence in lower-dimensional (e.g. 2-choice) decisions. Our results extend the observations of Webb et al. [14], obtained as emergent properties of training neural networks on high-dimensional stimulus sets, to the framework of classical SDT. An advantage of this approach is that we are able to precisely isolate a source of the detection-like confidence effects in higher-dimensional spaces in the normalisation induced by other (distractor) stimulus dimensions on confidence.
Previous work has also sought to link the emergence of a positive evidence bias to the degree of normalisation in the computation of confidence. Maniscalco et al. [12] proposed that the degree to which units within a neural network exhibit mutual inhibition predicts whether they will show discrimination- or detection-like confidence profiles. Less inhibition tuning led to detection-like confidence profiles, and accounted for empirical features of a positive evidence bias. This occurs because lower tuned inhibition leads to confidence being driven more by the absolute evidence than the difference in evidence between response options. In our SDT model, in contrast, detection-like confidence profiles are associated with greater normalisation of confidence within a higher-dimensional feature space. These two perspectives can be reconciled by recognising that in [12], the degree of inhibition is an intrinsic property of the network’s units, rather than being a consequence of a confidence computation. It is possible, then, that a detection-like “neuron” that shows lower tuned inhibition within a neural network is in fact exactly what would be required to approximate the more aggressive normalisation needed for computing (discrimination) confidence in higher-dimensional feature spaces.
Our model suggests that one source of unequal variance in signal detection models may be in the confidence computation itself, rather than (or in addition to) in the representation of evidence. This insight is also consistent with recent findings that the positive evidence bias (and associated variance aysmmetry) can be “flipped” just by changing the target dimension, with a positive evidence bias becoming a negative evidence bias if subjects are asked to report the lower value item in a pair [18]. Such a flip would be more difficult to accommodate if a positive evidence bias was driven by lower level asymmetries in the representation of sensory evidence and/or a function of (signed) input magnitude [7]. Similar asymmetrical distributions have been observed in the latent space of neural network representations of the stimulus input (obtained by using a variational autoencoder to extract a low-dimensional latent space underlying the high-dimensional stimulus inputs used in their experiments; [14]). However, while these models capture unequal variance as a consequence of features of the training distribution, they presumably would not be able to accommodate an instantaneous flip in the PEB without retraining the model [18]. Here, in contrast, we suggest a key driver of the sign of the PEB is the rapid and flexible influence of changing the goal or hypothesis space being entertained by the subject.
More generally, our results suggest, together with [14], that metacognitive biases may emerge due to subjects adopting a richer hypothesis space than assumed by the experimenter. This proposal aligns with other work showing how a mismatch between the experimenter’s and subjects’ task state-representations explains seemingly maladaptive behaviour in decision making [19] and metacognition [20, 21]. This hypothesis may be testable in future experiments that vary subjects’ expectations about the task-relevance of different stimuli. For instance, it would be of interest to know whether a positive evidence bias becomes stronger when a subject is induced to expect a richer stimulus space, even if the range of stimuli in play remain constant. These and other hypotheses that emerge from extending SDT into higher dimensions make for a rich source of hypotheses for future experiments.
A broader implication of our results is that, in higher dimensional spaces, decision rules for detection and identification become more similar [8, 10]. In line with this idea, Mazor et al. [8] showed that subjects incorporated detection-relevant evidence into their discrimination confidence, and discrimination-relevant evidence into their detection confidence. It is interesting to speculate that strict dissociations between detection and discrimination performance may be artefacts of the adoption of narrower stimulus spaces. Given that such dissociations have been foundational in establishing an empirical basis for unconscious perception (e.g. in blindsight) [22, 23, 24, 15, 25], establishing whether these findings generalise to more naturalistic stimulus spaces represents an important future direction for this research program.
Acknowledgments
This work was supported by a UKRI/EPSRC Programme Grant [EP/V000748/1]. SMF is a CIFAR Fellow in the Brain, Mind & Consciousness Program, and supported by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (selected as ERC Consolidator, grant number 101043666). Parts of this work were previously presented at the Cognitive Computational Neuroscience Meeting, Oxford, August 2023.
Appendix A Confidence can be computed using the softmax
As described in the Introduction, we can use Bayes’ rule to obtain a posterior distribution over stimulus categories given the sensory evidence $X$. Here we show that, assuming flat priors over stimulus categories, this reduces to the softmax of the evidence. We start by rewriting Equation 2:
$confidence=P(s_{a}|X)=\frac{p(X|s_{a})}{\sum_{i=1}^{k}p(X|s_{i})}=\frac{1}{\sum_{i=1}^{k}\frac{p(X|s_{i})}{p(X|s_{a})}}$ (5)
For any stimulus $s$, $p(X|s)$ is the likelihood of the joint normal distribution over $X$ with mean $\mu_{s}\in\{0,\mu_{target}\}^{k}$ and covariance $\Sigma=I_{k}$, where $\mu_{s}$ is a $k$-dimensional vector with $\mu_{target}$ on the $s$th coordinate and 0s elsewhere, $\mu_{target}$ is the observers’ sensitivity to the target dimension, and $I_{k}$ indicates the $k\times k$ identity matrix. The density of the multivariate normal distribution is defined as:
$p(X|\mu_{s},\Sigma=I_{k})$ $=\frac{1}{\sqrt{(2\pi)^{k}|\Sigma|}}e^{-\frac{1}{2}(X-\mu_{s})^{\top}\Sigma^{-1}(X-\mu_{s})}$ (6)
$=\frac{1}{\sqrt{(2\pi)^{k}}}e^{-\frac{1}{2}\|X-\mu_{s}\|^{2}}$ (7)
We can plug this into Equation 5 and simplify to derive confidence:
$confidence=P(s_{a}|X)$ $=\frac{1}{\sum_{i=1}^{k}\frac{p(X|s_{i})}{p(X|s_{a})}}$ (8)
$=\frac{1}{\sum_{i=1}^{k}\frac{e^{-\frac{1}{2}\|X-\mu_{s_{i}}\|^{2}}}{e^{-\frac{1}{2}\|X-\mu_{s_{a}}\|^{2}}}}$ (9)
$=\frac{1}{\sum_{i=1}^{k}e^{-\frac{1}{2}\left[\|X-\mu_{s_{i}}\|^{2}-\|X-\mu_{s_{a}}\|^{2}\right]}}$ (10)
$=\frac{1}{\sum_{i=1}^{k}e^{-\frac{1}{2}\left[(X_{i}-\mu_{target})^{2}+\sum_{j\neq i}X_{j}^{2}-(X_{a}-\mu_{target})^{2}-\sum_{j\neq a}X_{j}^{2}\right]}}$ (11)
$=\frac{1}{\sum_{i}e^{-\frac{1}{2}\left[(X_{i}-\mu_{target})^{2}+X_{a}^{2}+\sum_{j\notin\{a,i\}}X_{j}^{2}-(X_{a}-\mu_{target})^{2}-X_{i}^{2}-\sum_{j\notin\{a,i\}}X_{j}^{2}\right]}}$ (12)
$=\frac{1}{\sum_{i}e^{\mu_{target}(X_{i}-X_{a})}}$ (13)
$=\frac{e^{\mu_{target}X_{a}}}{\sum_{i}e^{\mu_{target}X_{i}}}$ (14)
$=\sigma_{\mu_{target}}(X)_{a}$ (15)
As a result, the posterior distribution $P(S|X)$ can be efficiently calculated using the softmax with inverse temperature $\beta=\mu_{target}$.
Appendix B The positive evidence bias grows with dimensionality
The positive evidence bias is defined as an increased sensitivity in confidence to the evidence for the chosen alternative, relative to evidence for unchosen alternatives. To quantify the positive evidence bias, then, we need to take the partial derivative of confidence with respect to each of the evidence dimensions $X_{i}$:
$\frac{\partial}{\partial X_{i}}P(s_{a}|X)$ $=\frac{\partial}{\partial X_{i}}\sigma_{\mu_{target}}(X)_{a}$ (16)
$=\frac{\partial}{\partial X_{i}}\frac{e^{\mu_{target}\,X_{a}}}{\sum_{j}e^{\mu_{target}\,X_{j}}}$ (17)
$=\frac{\frac{\partial}{\partial X_{i}}\left[e^{\mu_{target}\,X_{a}}\right]\sum_{j}e^{\mu_{target}\,X_{j}}-e^{\mu_{target}\,X_{a}}\frac{\partial}{\partial X_{i}}\left[\sum_{j}e^{\mu_{target}\,X_{j}}\right]}{(\sum_{j}e^{\mu_{target}\,X_{j}})^{2}}$ (18)
For $i=a$, this reduces to:
$\frac{\partial}{\partial X_{i}}P(s_{a}|X)$ $=\frac{\mu_{target}e^{\mu_{target}\,X_{a}}\sum_{j}e^{\mu_{target}\,X_{j}}-\mu_{target}e^{\mu_{target}\,X_{a}}e^{\mu_{target}\,X_{a}}}{(\sum_{j}e^{\mu_{target}\,X_{j}})^{2}}$
$=\mu_{target}\frac{e^{\mu_{target}\,X_{a}}}{\sum_{j}e^{\mu_{target}\,X_{j}}}\frac{\sum_{j}e^{\mu_{target}\,X_{j}}-e^{\mu_{target}\,X_{a}}}{\sum_{j}e^{\mu_{target}\,X_{j}}}$
$=\mu_{target}\sigma_{\mu_{target}}(X)_{a}(1-\sigma_{\mu_{target}}(X)_{a})$
For $i\neq a$, it reduces to:
$\frac{\partial}{\partial X_{i}}P(s_{a}|X)$ $=\frac{-\mu_{target}e^{\mu_{target}\,X_{a}}e^{\mu_{target}\,X_{i}}}{(\sum_{j}e^{\mu_{target}\,X_{j}})^{2}}$
$=-\mu_{target}\sigma_{\mu_{target}}(X)_{a}\sigma_{\mu_{target}}(X)_{i}$
Together, these define the Jacobian matrix $\mathbf{J}_{P(S|X)}=\mu_{target}\left[\textrm{diag}(\sigma_{\mu_{target}}(X))-\sigma_{\mu_{target}}(X)\sigma_{\mu_{target}}(X)^{\top}\right]$. Since the PEB occurs when the absolute change in confidence is greater with respect to $X_{a}$ than $X_{i}$ for $i\neq a$, we can determine this using the log ratio of the two partial derivatives:
$\textrm{PEB}_{i}=\textrm{log}\!\left\lvert\frac{\frac{\partial}{\partial X_{a}}P(s_{a}|X)}{\frac{\partial}{\partial X_{i}}P(s_{a}|X)}\right\rvert$ $=\textrm{log}\!\left(\frac{\mu_{target}\sigma_{\mu_{target}}(X)_{a}(1-\sigma_{\mu_{target}}(X)_{a})}{\mu_{target}\sigma_{\mu_{target}}(X)_{a}\sigma_{\mu_{target}}(X)_{i}}\right)$ (19)
$=\textrm{log}\!\left(\frac{1-\sigma_{\mu_{target}}(X)_{a}}{\sigma_{\mu_{target}}(X)_{i}}\right)$ (20)
$=\textrm{log}\!\left(\frac{1-\frac{e^{\mu_{target}\,X_{a}}}{\sum_{j}e^{\mu_{target}\,X_{j}}}}{\frac{e^{\mu_{target}\,X_{i}}}{\sum_{j}e^{\mu_{target}\,X_{j}}}}\right)$ (21)
$=\textrm{log}\!\left(\frac{\sum_{j}e^{\mu_{target}\,X_{j}}-e^{\mu_{target}\,X_{a}}}{e^{\mu_{target}\,X_{i}}}\right)$ (22)
$=\textrm{log}\!\left(\sum_{j}e^{\mu_{target}(X_{j}-X_{i})}-e^{\mu_{target}(X_{a}-X_{i})}\right)$ (23)
$=\textrm{log}\!\left(1+\sum_{j\notin\{a,i\}}e^{\mu_{target}(X_{j}-X_{i})}\right)$ (24)
When $\textrm{PEB}_{i}=0$, evidence for the chosen and unchosen options exert the same influence on confidence. The positive evidence bias arises when $\textrm{PEB}_{i}>0$, since in this case the evidence for the chosen option exerts a larger influence than the evidence for an unchosen option. $\textrm{PEB}_{i}=0$ for $k=2$ since there are no unchosen alternatives other than $i$, and so there are no terms in the sum. For $k>2$, $\textrm{PEB}_{i}>0$ since each additional term in the sum must be positive. If we take the limit as $k\rightarrow\infty$, we obtain:
$\lim_{k\to\infty}\textrm{PEB}_{i}=\lim_{k\to\infty}\textrm{log}\!\left(1+\sum_{j\notin\{a,i\}}e^{\mu_{target}(X_{j}-X_{i})}\right)=\infty$ (25)
So, for sufficiently large $k$, the influence of evidence for the chosen option on confidence will dwarf the influence of evidence for unchosen options, leading to detection-like confidence criteria.